# The State of Open Science in Software Engineering Research: A Case Study of ICSE Artifacts

### v1

@@ IV-B Answering RQ2: Types of Modifications
Figure 5: Types of Modifications
IV-C Answering RQ3: Challenges Preventing Executability
@@
Figure 6: List of Challenges
IV-D Answering RQ4: Reproducibility and Contributing Factors
@@
Figure 7: Reproducibility Status of Executable Artifacts
@@ V Discussion
V-A Actionable Guidelines for Improving Executability
@@
V-B Preparation and Evaluation of Our Replication Package
Preparation of our Replication Package. Our replication package was constructed following the actionable guidelines derived from executing and reproducing 100 ICSE replication packages in this study. It is organized into two complementary parts. Part 1 contains all analysis scripts and datasets required to regenerate the results reported in this paper. It was developed in full compliance with the documentation and packaging guidelines introduced in Section V-A. Along with the source code, we also provide a Dockerfile to run our experiments in an isolated environment. Part 2 provides a curated, per-artifact record of our execution and reproduction attempts for 100 ICSE replication packages. It does not modify the original artifacts but documents our interactions with them to ensure full traceability and transparency. Each artifact record includes four files (README.md, execution-log.md, patch.md, and updated-instructions.md) and one folder (/errors) when applicable. The detail purpose of these files can be found in the README.md file of our replication package [ReplicationPackage].
@@ V-B Preparation and Evaluation of Our Replication Package
$\bullet$ Assessment Summary (Part Two). Table III summarizes the assessment of Part II by the four evaluators (individual responses are available in our replication package [ReplicationPackage]). We evaluated this part using a stratified random sample of 20 artifacts drawn proportionally from the three executability categories (8 Executable, 5 Partially Executable, and 7 Not Executable) and distributed evenly among four evaluators. Each evaluator first reviewed the README.md file to verify the factual accuracy of our documentation. They then executed each artifact twice: once strictly following the original instructions and once using our meta-documentation (execution-log.md, patch.md, and updated-instructions.md). This two-pass execution allowed evaluators to assess both the baseline usability of the artifacts and the effectiveness of our structured documentation in improving comprehensibility and execution success.
TABLE III: External Evaluation Results for Part Two (E = Executable, count = 8; PE = Partially Executable, count = 5; NE = Not Executable, count = 7; T = Total, count = 20)
@@
Although we could not directly apply our proposed guidelines (Section V-A) to third-party artifacts, we compensated through structured meta-documentation that improved traceability and reproducibility. In line with G1, we evaluated each artifact against the five documentation components and filled missing details in our meta-files, which evaluators later confirmed as factually accurate. For G2, since we could not control environment leakage in external packages, we mitigated it by recording complete execution logs, patches, and errors and providing them in our replication package in a structured, hyperlinked format. Regarding G3, re-containerizing all 100 heterogeneous artifacts was infeasible; instead, we documented key environmental and configuration details, which evaluators reported as improving clarity and reusability except in cases where the original packages lacked sufficient information for full replication.
@@ VI Threats to Validity
Threats to construct validity concern the correctness and consistency of how we measured executability, effort, and reproducibility. Although these constructs are grounded in prior literature, some steps, such as identifying modification types or diagnosing failure causes, require judgment. We mitigated subjectivity bias through structured coding schemes, predefined criteria, and agreement across evaluators. When one evaluator could not execute an artifact, a second evaluator repeated the process independently. If both encountered the same issue, we attributed the failure to the artifact. Our effort estimates, based on agile-style bucketed time ranges, may be sensitive to changes in the boundaries. A sensitivity analysis showed that increasing the ranges by 5-10 percent produced no changes, while decreasing them caused only minor shifts, such as a few low-effort cases moving to moderate, indicating a negligible impact on overall findings. Hardware constraints present another construct validity threat. Some artifacts required extremely high-end setups, such as machines with 250 GB of RAM or multiple NVIDIA A100 GPUs, which we could not precisely replicate. We mitigated this limitation by using the closest available hardware (Section III-E) and adjusting permitted configurations, such as batch size and dependency versions, accordingly.
VII Related Work
@@
Artifact Sharing and Availability. Studies emphasize that open artifact sharing is essential. Oliveira Jr. et al. [OliveiraJr2024] and Cordeiro and Oliveira Jr. [Cordeiro2025a, Cordeiro2025b] advocate SE-specific frameworks and show that artifact badges improve availability. Liu et al. [Liu2024] confirm a rising trend in sharing but warn that availability alone does not ensure usability. Others highlight reproducible research as a standard practice [Madeyski2017] and call for structured evaluations, especially in MSR [Tkachenko2025, GonzalezBarahona2023].
Executability Challenges. Despite wider sharing, executability remains difficult. Mukhtar et al. [Mukhtar2024] found missing dependencies and poor documentation in SE tools. Ajayi et al. [Ajayi2023] and Bonneel et al. [Bonneel2020] report failures due to configuration drift. Proposed solutions include containerization [Costa2024] and modeling deep environment variability [Acher2024].
Reproducibility Audits. Several audits show that execution does not imply reproducibility. Nong et al. [Nong2023], Agal et al. [Agal2025], and Liu et al. [Liu2022] cite inconsistent pipelines and missing criteria. Earlier, Mahmood et al. [Mahmood2018] and Rodriguez-Perez et al. [RodriguezPerez2018] highlighted gaps in defect prediction reproducibility. Robles and Gonzalez-Barahona [GonzalezBarahona2023] proposed structured MSR assessment protocols.
Community Standards and Taxonomies. Initiatives like the ACM artifact badge clarify reproducibility terms. Gundersen and Kjensmo [Gundersen2018] introduced a taxonomy; Cordeiro and Oliveira Jr. [Cordeiro2025a] emphasized usability and clarity. Acher et al. [Acher2024] stress modeling software variability for reproducibility.
@@
VIII Conclusion
Despite the growing emphasis on open science in SE research, the practical usability of shared artifacts remains insufficiently examined. This study provides the first large-scale empirical assessment of executability and reproducibility across 100 ICSE replication packages. By investing roughly 650 person-hours of analysis, we uncovered substantial gaps: only 40% of artifacts were executable, just one-third ran without modification, and reproducibility was achieved in only 35% of those that executed. To help close this gap, we proposed three actionable guidelines: adopt comprehensive documentation, validate artifacts in clean environments to avoid configuration leakage, and supply both containerized setups and accessible source code (where applicable). Our findings show that artifacts including these components achieve substantially higher executability and reproducibility. Strengthening open science in SE requires not only sharing artifacts but ensuring they are runnable, verifiable, and usable by others. Following the guidelines introduced in this study can help researchers and conference organizers advance reproducible and trustworthy SE research.
IX Acknowledgment

### v2

@@ 4.2. Answering RQ2: Types of Modifications Needed to Execute Replication Packages
Figure 6. List of Challenges
A hierarchical chart displaying execution barriers organized into 3 main themes: Documentation Issues, Environment Setup Challenges, and Technical Issues. Each theme branches into specific challenge types with frequency counts.
4.3. Answering RQ3: Challenges that Prevent Executability of Replication Packages
@@
4.4. Answering RQ4: Reproducibility Outcomes and Contributing Factors
@@
Figure 7. Reproducibility Status of Executable Artifacts
A stacked bar chart showing reproducibility results for 64 executable artifacts. The chart is divided into three colored segments: fully reproducible (14 artifacts, 14%), partially reproducible (16 artifacts, 16%), and not reproducible (34 artifacts, 70%).
@@ 5. Discussion
5.1. Actionable Guidelines for Improving Executability
@@
5.2. Preparation and Evaluation of Our Replication Package
Preparation of our Replication Package. Our replication package was constructed following the actionable guidelines derived from executing and reproducing 100 ICSE replication packages in this study. It is organized into two complementary parts. Part 1 contains all analysis scripts and datasets required to regenerate the results reported in this paper. It was developed in full compliance with the documentation and packaging guidelines introduced in Section 5.1. Along with the source code, we also provide a Dockerfile to run our experiments in an isolated environment. Part 2 provides a curated, per-artifact record of our execution and reproduction attempts for 100 ICSE replication packages. It does not modify the original artifacts but documents our interactions with them to ensure full traceability and transparency. Each artifact record includes four files (README.md, execution-log.md, patch.md, and updated-instructions.md) and one folder (/errors) when applicable. The detail purpose of these files can be found in the README.md file of our replication package (Replication Package, 2026).
@@ 5.2. Preparation and Evaluation of Our Replication Package
$\bullet$ Assessment Summary (Part Two). Table 3 summarizes the assessment of Part II by the four evaluators (individual responses are available in our replication package (Replication Package, 2026)). We evaluated this part using a stratified random sample of 20 artifacts drawn proportionally from the three executability categories (8 Executable, 5 Partially Executable, and 7 Not Executable) and distributed evenly among four evaluators. Each evaluator first reviewed the README.md file to verify the factual accuracy of our documentation. They then executed each artifact twice: once strictly following the original instructions and once using our meta-documentation (execution-log.md, patch.md, and updated-instructions.md). This two-pass execution allowed evaluators to assess both the baseline usability of the artifacts and the effectiveness of our structured documentation in improving comprehensibility and execution success.
Table 3. External Evaluation Results for Part Two (E = Executable, count = 8; PE = Partially Executable, count = 5; NE = Not Executable, count = 7; T = Total, count = 20)
@@
Although we could not directly apply our proposed guidelines (Section 5.1) to third-party artifacts, we compensated through structured meta-documentation that improved traceability and reproducibility. In line with G1, we evaluated each artifact against the five documentation components and filled missing details in our meta-files, which evaluators later confirmed as factually accurate. For G2, since we could not control environment leakage in external packages, we mitigated it by recording complete execution logs, patches, and errors and providing them in our replication package in a structured, hyperlinked format. Regarding G3, re-containerizing all 100 heterogeneous artifacts was infeasible; instead, we documented key environmental and configuration details, which evaluators reported as improving clarity and reusability except in cases where the original packages lacked sufficient information for full replication.
@@ 6. Threats to Validity
Threats to construct validity concern the correctness and consistency of how we measured executability, effort, and reproducibility. Although these constructs are grounded in prior literature, some steps, such as identifying modification types or diagnosing failure causes, require judgment. We mitigated subjectivity bias through structured coding schemes, predefined criteria, and agreement across evaluators. When one evaluator could not execute an artifact, a second evaluator repeated the process independently. If both encountered the same issue, we attributed the failure to the artifact. Our effort estimates, based on agile-style bucketed time ranges, may be sensitive to changes in the boundaries. A sensitivity analysis showed that increasing the ranges by 5-10 percent produced no changes, while decreasing them caused only minor shifts, such as a few low-effort cases moving to moderate, indicating a negligible impact on overall findings. Hardware constraints present another construct validity threat. Some artifacts required extremely high-end setups, such as machines with 250 GB of RAM or multiple NVIDIA A100 GPUs, which we could not precisely replicate. We mitigated this limitation by using the closest available hardware (Section 3.5) and adjusting permitted configurations, such as batch size and dependency versions, accordingly.
7. Related Work
@@
Artifact Sharing and Availability. Studies emphasize that open artifact sharing is essential. Oliveira Jr. et al. (Oliveira Jr et al., 2024) and Cordeiro and Oliveira Jr. (Cordeiro and Oliveira Jr, 2025a, b) advocate SE-specific frameworks and show that artifact badges improve availability. Liu et al. (Liu et al., 2024) confirm a rising trend in sharing but warn that availability alone does not ensure usability. Others highlight reproducible research as a standard practice (Madeyski and Kitchenham, 2017) and call for structured evaluations, especially in MSR (Tkachenko, 2025; Gonzalez-Barahona and Robles, 2023).
Executability Challenges. Despite wider sharing, executability remains difficult. Mukhtar et al. (Mukhtar et al., 2024) found missing dependencies and poor documentation in SE tools. Ajayi et al. (Ajayi et al., 2023) and Bonneel et al. (Bonneel et al., 2020) report failures due to configuration drift. Proposed solutions include containerization (Costa et al., 2024) and modeling deep environment variability (Acher et al., 2024).
Reproducibility Audits. Several audits show that execution does not imply reproducibility. Nong et al. (Nong et al., 2023), Agal et al. (Agal et al., 2025), and Liu et al. (Liu et al., 2022) cite inconsistent pipelines and missing criteria. Earlier, Mahmood et al. (Mahmood et al., 2018) and Rodriguez-Perez et al. (Rodríguez-Pérez et al., 2018a) highlighted gaps in defect prediction reproducibility. Robles and Gonzalez-Barahona (Gonzalez-Barahona and Robles, 2023) proposed structured MSR assessment protocols.
Community Standards and Taxonomies. Initiatives like the ACM artifact badge clarify reproducibility terms. Gundersen and Kjensmo (Gundersen and Kjensmo, 2018a) introduced a taxonomy; Cordeiro and Oliveira Jr. (Cordeiro and Oliveira Jr, 2025a) emphasized usability and clarity. Acher et al. (Acher et al., 2024) stress modeling software variability for reproducibility.
@@
8. Conclusion
Despite the growing emphasis on open science in SE research, the practical usability of shared artifacts remains insufficiently examined. This study provides the first large-scale empirical assessment of executability and reproducibility across 100 ICSE replication packages. By investing roughly 650 person-hours of analysis, we uncovered substantial gaps; only 40% of artifacts were executable, just one-third ran without modification, and reproducibility was achieved in only 35% of those that executed. To help close this gap, we proposed three actionable guidelines; adopt comprehensive documentation, validate artifacts in clean environments to avoid configuration leakage, and supply both containerized setups and accessible source code (where applicable). Our findings show that artifacts including these components achieve substantially higher executability and reproducibility. Strengthening open science in SE requires not only sharing artifacts but ensuring they are runnable, verifiable, and usable by others. Following the guidelines introduced in this study can help researchers and conference organizers advance reproducible and trustworthy SE research.
Acknowledgements.